OpenStudy (anonymous):

I have to use the quadratic formula for 3x^2 + 14x = 8. I have already done the problem in class however I am confused on the last couple of steps.

OpenStudy (luigi0210):

What coupe of steps?

OpenStudy (anonymous):

\[\frac{ -14 \pm10 }{ 6 }\] Thats where I am stuck

OpenStudy (luigi0210):

Break it up into two fractions.

OpenStudy (oaktree):

Okay. So we split into two different parts. Let's take the plus first. We have\[\frac{ -14 + 10 }{ 6 } = \frac{-4}{6}\]Now you need to simplify that. Do you know how?

OpenStudy (anonymous):

divide by two?

OpenStudy (luigi0210):

Make one a plus and another a minus

OpenStudy (anonymous):

@OakTree am I right :)

OpenStudy (anonymous):

omg yes I finished the rest of it :D